For the reaction ,

2NO₂ (ɡ) ⇌N₂O₄ (ɡ)
Kp is 7.14× 10⁻⁵ Pa at 303 K . Find Kc at this temperature.
1
Expert's answer
2019-10-31T07:29:04-0400

Reaction is:


"2NO_2\to N_2O_4"

Here,

"\Delta n=1-2=-1\\also,\\"


"k_p=k_c\times(RT)^{\Delta n}\\or\\k_c=\frac{k_p}{(RT)^{-1}}=k_p(RT)=7.14\times10^{-5}\times8.314\times303\\k_c=0.17986\ Pa\ J\ mol^{-1}"